OpenStudy (anonymous):

Please help!! Find the nth partial sum of the arithmetic sequence: 1.50, 1.45, 1.40, 1.35 . . . n=20

OpenStudy (anonymous):

I think it is 0.55, is that correct?

OpenStudy (cwrw238):

not sure what you mean by nth partial sum

jimthompson5910 (jim_thompson5910):

It just means add up the first n terms. In this case, n = 20, so add up the first 20 terms.

OpenStudy (anonymous):

ok Is 0.55 correct, can you just tell me?

jimthompson5910 (jim_thompson5910):

no, you have 20 terms which most are over 0.55, so it's definitely not possible to add positive numbers over 0.55 to get 0.55

OpenStudy (cwrw238):

sum of 20 terms = 10(2*1.5 + (20-1)*-0.5)

OpenStudy (anonymous):

can you show me what to do? and how to get the answer?

OpenStudy (cwrw238):

* 10(2*1.5 + 19*-0.05) difference -0.05 not -0.5

jimthompson5910 (jim_thompson5910):

Sum of first n terms = (number of terms)*(first term+nth term)/2

jimthompson5910 (jim_thompson5910):

So S = (n*(a1 + an))/2

OpenStudy (anonymous):

r=1.45-1.50=-0.05 \[u _{1}=1.50\rightarrow u _{n}=1.55-n0.05\]\[\sum_{n=1}^{20}u _{n}=20\times1.55-0.05\times20\times21\div2=20.5\]Or your question is\[u _{20}=20\times(-0.05)+1.55=1.55-1=0.55\]

OpenStudy (cwrw238):

formula for sum n terms = (n/2)(2a + (n-1)d)

OpenStudy (anonymous):

So my answer, 0.55, is right then syderitic?

OpenStudy (anonymous):

@cwrw238 is that right?

jimthompson5910 (jim_thompson5910):

no, the answer isn't 0.55

OpenStudy (anonymous):

then idk what is?

OpenStudy (anonymous):

look, if you want the sum of the first 20 terms, then the answer is 20.5 if you want the 20th term of that sequence then the answer is 0.55

jimthompson5910 (jim_thompson5910):

Plug n = 20, a1 = 1.5 and an = 0.55 into S = (n*(a1 + an))/2 to find the sum

OpenStudy (anonymous):

it says Find the nth partial sum @Syderitic

OpenStudy (anonymous):

So @jim_thompson5910 I got 20.5. are u sure that is it?

OpenStudy (anonymous):

English is not my primary language... sry

OpenStudy (anonymous):

i would go for the 0.55 since you want the partial sum, so i'm guessing that's a term.

jimthompson5910 (jim_thompson5910):

you got the right answer, I'm getting 20.5 as well

OpenStudy (anonymous):

Ok ughhh idk if i should put .55 or 20.5

OpenStudy (anonymous):

because they want the nth PARTIAL sum

jimthompson5910 (jim_thompson5910):

0.55 is the 20th term, there's no way it's the 20th partial sum

OpenStudy (cwrw238):

i make it 20.5

OpenStudy (anonymous):

are u guys positive?

jimthompson5910 (jim_thompson5910):

it's 20.5

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

the only thing i'm not positive of is the question itself...

OpenStudy (anonymous):

i am putting 20.5

OpenStudy (cwrw238):

partial sum is confusing

jimthompson5910 (jim_thompson5910):

If you really really want, you can add up the terms 1.5+1.45+1.4+1.35+1.3+1.25+1.2+1.15+1.1+1.05+1+0.95+0.9+0.85+0.8+0.75+0.7+0.65+0.6+0.55 These are the first 20 terms And you'll get 20.5 So this confirms the answer.

OpenStudy (anonymous):

thx